# LLaNA: Large Language and NeRF Assistant

### v1

@@ Appendix D Ground truth images and point clouds
Table 8: NeRF captioning - brief descriptions on ShapeNeRF–Text dataset. Baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
@@
LLaVA-vicuna-13b Image (Front View) 68.32 67.35 27.46 26.62 24.40
@@
Table 9: NeRF captioning - detailed descriptions on ShapeNeRF–Text dataset. Baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
@@
LLaVA-vicuna-13b Image (Front View) 78.40 75.68 22.65 33.04 35.70
@@
Table 10: NeRF Q&A - single-round Q&A on ShapeNeRF–Text dataset. Baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
Appendix E Additional qualitative examples
This section provides additional qualitative comparisons between the proposed method, i.e. LLaNA which directly processes NeRF, and the baselines that take as input images [44, 39] or 3D representations [77, 58]. In particular, Figs. 9, 10 and 11 show additional brief descriptions, detailed descriptions, and single-round Q&A provided as output by the different methods. Many examples, such as the white speaker in the third row of Fig. 10, are not described properly by MLLMs operating on point clouds. Indeed, due to the input point cloud containing only $8192$ points, these methods cannot perceive the object details, such as the curved surface of the speaker, therefore they predict that the object is a “cubic white object” or an “ice cube”. In other examples, such as the white screen sample in the last row of Fig. 11, the LLM operating on images cannot give the right answer to the question on the button location as it is not visible from the given viewpoint. Contrarily, by operating directly on the holistic representation provided by NeRFs, LLaNA provides the right answer in these situations.
@@ Appendix E Additional qualitative examples
Appendix F Information about datasets, models, and source code licenses

### v2

@@ Appendix E Ground-truth images and point clouds
Table 14: NeRF brief captioning on ShapeNeRF–Text dataset. Frozen baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
(FV: front-view)
@@
LLaVA-vicuna-13b Image (FV) 68.32 67.35 27.46 26.62 24.40
@@
Table 15: NeRF detailed captioning on ShapeNeRF–Text dataset. Frozen baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
@@
LLaVA-vicuna-13b Image (FV) 78.40 75.68 22.65 33.04 35.70
@@
Table 16: NeRF single-round Q&A on ShapeNeRF–Text dataset. Frozen baseline results obtained on data extracted from ShapeNet mesh data. Best results in bold. Runner-up underlined.
Appendix F Generalization experiments
We conducted an experiment to probe the generalization capabilities of LLaNA against the trained baselines. We evaluate the models on the subset of 200 Objaverse [13] objects with human-annotated captions used as a test set by PointLLM [78]. This evaluation protocol sets forth a challenging out-of-domain and open-set experiment (164 out of 200 Objaverse objects belong to categories not present in ShapeNeRF-Text). To test LLaNA, we fit NeRFs for all the objects of the test set. Then, we extracted colored point clouds and rendered front views from NeRFs, in order to test the baselines. In Tab. 17 we can observe that the scores of all models are significantly lower compared to Tab. 7, which hints at all models struggling when evaluated on objects very different from those belonging to the training domain. LLaNA achieves the second-best generalization performance after PointLLM. Yet, it is worth highlighting that the frozen modality-specific encoder of PointLLM (and GPT4Point) is PointBERT, which was pre-trained on Objaverse. In contrast, LLaNA meta-encoder, nf2vec, has been trained only on ShapeNet, meaning it has never encountered objects outside the ShapeNet categories.
Table 17: Generalization results on Objaverse.
NeRF captioning
Model Modality Sentence-BERT SimCSE BLEU-1 ROUGE-L METEOR
LLaVA-vicuna-13b Image (FV) 27.07 26.82 4.41 6.81 9.77
PointLLM-7b Point cloud 33.88 33.04 5.37 8.14 12.28
GPT4Point-Opt-2.7b Point cloud 25.94 29.04 4.25 7.99 10.42
LLaNA-7b NeRF 30.07 28.17 4.93 7.47 11.15
Appendix G Additional qualitative examples
This section provides additional qualitative comparisons between the proposed method, i.e. LLaNA which directly processes NeRF, and the baselines that take as input images [44, 39] or 3D representations [78, 58]. In particular, Figs. 12, 13 and 14 show additional brief descriptions, detailed descriptions, and single-round Q&A provided as output by the different methods. Many examples, such as the white speaker in the third row of Fig. 13, are not described properly by MLLMs operating on point clouds. Indeed, due to the input point cloud containing only $8192$ points, these methods cannot perceive the object details, such as the curved surface of the speaker, therefore they predict that the object is a “cubic white object” or an “ice cube”. In other examples, such as the white screen sample in the last row of Fig. 14, the LLM operating on images cannot give the right answer to the question on the button location as it is not visible from the given viewpoint. Contrarily, by operating directly on the holistic representation provided by NeRFs, LLaNA provides the right answer in these situations.
@@ Appendix G Additional qualitative examples
Appendix H Information about datasets, models, and source code licenses